OpenStudy (anonymous):

How do I expand: (x-i)(x+i)(x-4+i)(x-4-i)(x-2+i)(x-2-i)? Thanks in advance!

OpenStudy (badhi):

do you see (a-b)(a+b) pattern?

OpenStudy (anonymous):

You mean the (x-4-i)(x-4+i), etc.?

OpenStudy (badhi):

yeah.. and notice the property \((a+b)(a-b) = (a^2-b^2)\)

OpenStudy (anonymous):

Oh yeah, between the i's?

OpenStudy (anonymous):

(x-i)(x+i) = (x^2-i^2) = x^ + 1?

OpenStudy (anonymous):

x^2 + 1*

OpenStudy (badhi):

yes

OpenStudy (anonymous):

Forget it, I got it,

OpenStudy (zzr0ck3r):

\[\tiny (\hspace{0.2cm}x \hspace{0.2cm} - \hspace{0.2cm} i\hspace{0.2cm})\hspace{0.2cm}( \hspace{0.2cm}x\hspace{0.2cm}+\hspace{0.2cm}i\hspace{0.2cm})\hspace{0.2cm}(\hspace{0.2cm}x\hspace{0.2cm}-\hspace{0.2cm}4\hspace{0.2cm}+\hspace{0.2cm}i\hspace{0.2cm}\hspace{0.2cm})\hspace{0.2cm}(\hspace{0.2cm}x\hspace{0.2cm}-\hspace{0.2cm}4\hspace{0.2cm}-\hspace{0.2cm}i\hspace{0.2cm})\hspace{0.2cm}(\hspace{0.2cm}x\hspace{0.2cm}-\hspace{0.2cm}2\hspace{0.2cm}+\hspace{0.2cm}i\hspace{0.2cm})\hspace{0.2cm}(\hspace{0.2cm}\hspace{0.2cm}x\hspace{0.2cm}-\hspace{0.2cm}2\hspace{0.2cm}-\hspace{0.2cm}i\hspace{0.2cm})\hspace{0.2cm}\]

OpenStudy (zzr0ck3r):

Not what you meant?

OpenStudy (anonymous):

Yeah it was @zzr0ck3r, I was able to figure it out in the end, the question asked to form a polynomial function out of the given degree and zero values using real coefficients so I just really needed help expanding it more than anything. I managed to solve it though I believe if you check my work ^ Anyway, thanks @BAdhi and @zzr0ck3r